# The exact travelling wave solutions of a KPP equation

### v4

@@ Abstract
We obtain the exact analytical traveling wave solutions of the Kolmogorov-Petrovskii-Piskunov equation with the reaction term belonging to the class of functions, which includes that of the (generalized) Fisher equation, for the particular values of the waves speed. The solutions are written in terms of elementary functions.
@@ II Travelling waves
III The ODE which doesn’t contain explicitly the independent variable
Equation (5) doesn’t contain explicitly the independent variable $\xi$. This prompts the idea benguria ; kogan3 ; kogan4 ; kogan5 to consider $u$ as the new independent variable and
@@ III The ODE which doesn’t contain explicitly the independent variable
IV The correct boundary conditions for Eq. (9)
@@
After we understood the true nature of the boundary conditions (6) for (5), the situation with Eq. (9) becomes obvious. Note that the boundary conditions (10) are also exceptional goursat $p_{u}$ is not defined either at $u=0$ or at $u=1$.
However, looking back at (24) we realize that the boundary condition (10a) should be specified as follows: For $u\to 1$
@@
For practical purposes we can sometimes continue to use this boundary condition in the previous form (10a), like we deed in Section III. The differential equation itself produces the correct condition (30). However, if we want, for example, to solve (9) numerically, the condition (30) should be used from the beginning.
And looking back at our analysis of (19) we realize that any solution of Eq. (9) will satisfy (10b), so the second boundary condition (10b) can be discarded. It is satisfied automatically.
@@ Appendix C Perturbation theory II
In Section III we have solved Eq. (13) for $c=c_{0}$. Now let us study the solution for $c$ close to $c_{0}$, considering $\delta c\equiv c-c_{0}$ as the small parameter and using the perturbation series. To make life simpler $n=1$, we’ll consider only the case $n=1$. Presenting the solution as

### v5

@@ Abstract
We obtain the exact analytical traveling wave solutions of the Kolmogorov-Petrovskii-Piskunov equation with the reaction term belonging to the class of functions, which includes that of the (generalized) Fisher equation, for the particular values of the waves speed. Additionally we obtain the the exact analytical traveling wave solutions of the generalized Burgers–Huxley equation.
@@ I Introduction
Equation (5) is a nonlinear nonintegrable differential equation. The Cauchy problem cannot be solved by the inverse scattering transform for this equation and the problem of finding some exact solutions for this equation is an important task. Several methods of obtaining such solutions has been developped. We will mention here a few of them: the tanh-function method 4 ; 5 ; 6 , the Jacobi elliptic function expansion method 8 , the simplest equation method 11 ; 12 , the Exp-function method 14 , the $G^{\prime}/G$ expansion method 15 ; 16 ; 17 ; 18 , and the Kudryashov method 19 .
In this paper we will use the method of exclusion of the independent variable benguria ; kogan3 ; kogan4 ; kogan5 , which allows to reduce the problem of integration of the second order ODE to the problem of integration of the first order ODE, to obtain the exact solutions of Eq. (5). Thus we were able to obtain the exact solutions of the KPP equation which to the best of our knowledge were not known before. Additionally we obtain the the exact analytical traveling wave solutions of the generalized Burgers–Huxley equation.
II The ODE which doesn’t contain explicitly the independent variable
Equation (5) doesn’t contain explicitly the independent variable $\xi$. This prompts the idea to consider $u$ as the new independent variable and
@@ II The ODE which doesn’t contain explicitly the independent variable
III The correct boundary conditions for Eq. (9)
@@
After we understood the true nature of the boundary conditions (6) for (5), the situation with the boundary conditions (10) for Eq. (9) becomes obvious. Note that the boundary conditions (10) are also exceptional goursat $p_{u}$ is not defined either at $u=0$ or at $u=1$.
In the vicinity of $u=1$, Eq. (9) can be simplified to
$qq_{v}-cq=|F^{\prime}(1)|v,$ (30)
where again $v=1-u$, and $q\equiv dv/d\xi$. Equation (30) has two solutions
$q$ $=s_{1}v,$ (31a)
$q$ $=s_{2}v,$ (31b)
where $s_{1,2}$ are given by (23). Only the solution (31a) satisfies the condition (6a). Hence we realize that the boundary condition (10a) should be specified as follows: For $u\to 1$
@@
(Not the difference between the proportionality sign in (24) and the equality sign in (32).) One can easily check up that our solution (14) satisfies this boundary condition.
In the vicinity of $u=0$, Eq. (9) can be simplified to
$pp_{v}+cp=F^{\prime}(0)u.$ (33)
Equation (33) has two solutions
$p$ $=s_{3}u,$ (34a)
$p$ $=s_{4}u,$ (34b)
where $s_{3,4}$ are given by (26). Because both $s_{3}$ and $s_{4}$ have negative real part, both solutions will satisfy the boundary condition (6b). So we understood that the second boundary condition (10b) can be discarded – it is satisfied automatically.
@@ Appendix C Perturbation theory II
In Section II we have solved Eq. (13) for $c=c_{0}$. Now let us study the solution for $c$ close to $c_{0}$, considering $\delta c\equiv c-c_{0}$ as the small parameter and using the perturbation series. To make life simpler $n=1$, we’ll consider only the case $n=1$. Presenting the solution as